Suppose that a country has a tax rate on dividends (TP) of 40%. It has a tax rate on capital gains (TG) of 20%.

If these taxes affect the stock price on the ex-dividend date, what should happen to the stock price on the

ex-dividend date if a corporation pays a dividend per share of $5

a) A $10 decrease in the stock price.

b) A $3.75 decrease in the stock price.

c) A $3 decrease in the stock price.

d) A $2 decrease in the stock price.
Answer #1

(PO - PX) / D = (1 - tP) / (1-TG)

(PO - PX) /5 = ( 1-0.4) / ( 1-0.2)

PO - PX = 3.75

Option A $3.75 decrease in the stock price.

Remarks: PO is the stock price before ex and PX is the stock price after dividend.
Answer #2

To determine the impact of taxes on the stock price on the ex-dividend date, we need to consider the tax rates on dividends (TP) and capital gains (TG) and how they affect investors' after-tax returns.

On the ex-dividend date, when a corporation pays a dividend per share of $5, the stock price typically drops by the amount of the dividend. This is because investors who purchase the stock after the ex-dividend date are not entitled to receive the dividend payment, so the stock price adjusts to reflect the reduced value (the amount of the dividend is effectively taken out of the stock price).

Let's calculate the after-tax return for investors in each scenario:

a) A $10 decrease in the stock price:

After-tax return for dividends (TP = 40%): $5 * (1 - 0.40) = $3

After-tax return for capital gains (TG = 20%): $10 * (1 - 0.20) = $8

Total after-tax return for investors: $3 (dividends) + $8 (capital gains) = $11

b) A $3.75 decrease in the stock price:

After-tax return for dividends (TP = 40%): $5 * (1 - 0.40) = $3

After-tax return for capital gains (TG = 20%): $3.75 * (1 - 0.20) = $3

Total after-tax return for investors: $3 (dividends) + $3 (capital gains) = $6

c) A $3 decrease in the stock price:

After-tax return for dividends (TP = 40%): $5 * (1 - 0.40) = $3

After-tax return for capital gains (TG = 20%): $3 * (1 - 0.20) = $2.40

Total after-tax return for investors: $3 (dividends) + $2.40 (capital gains) = $5.40

d) A $2 decrease in the stock price:

After-tax return for dividends (TP = 40%): $5 * (1 - 0.40) = $3

After-tax return for capital gains (TG = 20%): $2 * (1 - 0.20) = $1.60

Total after-tax return for investors: $3 (dividends) + $1.60 (capital gains) = $4.60

Based on the calculations, option (a) with a $10 decrease in the stock price results in the highest total after-tax return for investors ($11). Therefore, the stock price should decrease by $10 on the ex-dividend date if a corporation pays a dividend per share of $5. Option (a) is the correct answer.


answered by: mervetokaz